The estimated supply function for avocados is Q = 58 + 15p - 20pf, where p is the price of avocados and pf is the price of fertilizer. When the price of fertilizer increases by $1.50 the inverse supply curve shifts The price elasticity of supply, estimated at p = 4 and pf = 10, is approximately When the price of fertilizer increases by $1.50 the supply curve shifts
Answer #1

As per the equation, Q = 58 + 15p - 20pf,

there is a direct relation between the price of avocados and the supply of avocados. It means that as the price of avocados increases, the demand of avocados also increases. There is an inverse relation between the price of fertilizers and the quantity supplied of avocados. It means that as the price of fertilizers increases, the quantity supplied of avocados reduces.

When the price of fertilizer increases, the supply curve shifts leftwards stating that the quantity supplied reduces.

Quantity supplied at p= 4, and pf = 10

Q = 58 + 15p - 20pf

Q = 58 + 15 *4 - 20*10

= 58 + 60 - 200

= -82

If the price of fertilizer increases by $1.50

Q = 58 + 15*4 - 20*11.50

= 58 + 60 - 230

= -112

So, Elasticity of supply = Percentage change in Quantity supplied / Percentage change in price

= AQ/Q1* P1/AP

= -112 - (-82) / -82 * 10 / 11.50-10

= -112+82 / -82 * 10 / 1.50

= -30 / -82 * 10/ 1.50

= 2.44

Thus it is highly elastic supply.